# DGP-LVM: Derivative Gaussian process latent variable models

### v2

@@ 4.1 Simulated data
We consider two scenarios to generate simulated data. In our first scenario, we generate data from a multi-output GP with scaled derivative SE covariance function with varying hyperparameters and correlated outputs (see Section 3). This constitutes the case where DGP-LVMs align with the underlying process. However, datasets generated this way can vary strongly in the amount of signal they contain, thus adding a lot of random variation in the simulation results. To account for this issue, in our second scenario, we generate data from a derivative periodic process with the true generating function
$f_{ij}=\alpha_{j}\sin\left(\frac{x_{i}}{\rho_{j}}\right)\quad\text{and}\quad f^{\prime}_{ij}=\frac{\alpha^{\prime}_{j}}{\rho_{j}}\cos\left(\frac{x_{i}}{\rho_{j}}\right)$ (16)
and corresponding data simulations
@@
The hyperparameters of this data generating process fulfill a similar purpose than those of SE derivative GPs, hence we choose to use the same hyperparameter names for simplicity. This second scenario is important on two counts. First, it allows us to better control the amount of signal contained in each generated dataset. Second, it demonstrates that DGP-LVM can achieve good results even when the underlying generating process in not actually a GP.
@@
To demonstrate the adversity of scale difference between $y$ and $y^{\prime}$, we induce a scaling factor of $\lambda=3$ that propagates through the GP marginal SD and error SD (see Section 3.1.1). The GP marginal SD for the output $y$ and the derivative $y^{\prime}$ are related through $\lambda$ such that $\alpha=\lambda\alpha^{\prime}$. Similarly for the error SD, $\sigma=\lambda\sigma^{\prime}$. Therefore, we only specify sampling distributions for $\alpha^{\prime}\sim\text{Uniform}(0.4,0.6)$ and $\sigma^{\prime}\sim\text{Uniform}(0.05,0.15)$. The ground truth GP length scale is sampled as $\rho\sim\text{Uniform}(0.5,1)$. The choice of our sampling distributions enabled us to explicitly select a range of values for our hyperparameters for which the simulated GP data would contain enough signals. In reality, $\lambda>3$ or $\lambda<1/3$ may very well occur (see Section 5). In the simulations, we chose to avoid more extreme $\lambda$ values to prevent substantial convergence issues for models without scaling modifications. This allows us to showcase these models’ (reduced) performance without confounding this finding with convergence considerations. Further, we introduce an uniform between-dimension correlation of $0.5$ for the simulated GP scenario as to represent moderate interactions between outputs. Combined with the sampling of true hyperparameters for each output dimension, our simulated data mimics the real-world data scenario where such assumptions are prevalent.
@@
Lastly, we generate ground truth for $x$ as a sequence of values between {0.5, 10} with equidistant steps of 0.5 and we choose a value for prior measurement SD of the noisy $\tilde{x}$ as $s=0.3$ (see Section 3.3). This resembles the realistic scenario where observed times are already a relatively good measure of latent pseudotime considering the overall input scale. Both simulation studies are generated as multi-output data with three sets of output dimensions namely $D=5,10\text{ }\text{and}\text{ }20$.
To keep model estimation times within manageable bounds, each simulated dataset contains only 20 observed and 20 derivative sample points. We perform 50 trials per simulation scenario, that is, generate 50 datasets for the GP and the periodic scenario, respectively.
@@ 4.2 Model setup
The models within our DGP-LVM framework can vary specifically in four components, namely the inclusion of (1) derivative information, (2) scaled derivatives, (3) varying hyperparameters, as well as (4) correlations across output dimensions. In order to study the individual importance of the four components, we systematically enable/disable each of them and investigate the resulting models’ performances. The underlying data generating process contains all of the above components, so any model that only has a subset of components will be misspecified at least to some degree. In our simulations, component combinations were fully crossed where sensible (see Table 1 for an overview). We fit 12 GP models for each selected number of output dimensions $D$ resulting in a total of 36 models fitted per generated dataset for a specific simulation scenario. We only exclude specific, non-sensible combinations. For example, it does not make sense to ask if a GP model, which does not include derivative information, accounts for the scaling of the derivatives. Due to substantial computation requirements of such simulation studies, we only considered models with SE covariance functions. Other Matern kernels (specifically Matern 3/2) are applied in the case study in Section 5.
@@
We use weakly-informative priors for all GP hyperparameters. For this purpose, we scale the priors of the marginal and error SD parameters ($\alpha$, $\sigma$ and their derivative counterparts $\alpha^{\prime}$, $\sigma^{\prime}$) by the empirical standard deviations of the data denoted by $s_{y}$ and $s_{y^{\prime}}$ such that they follow $\text{Normal}(0,s_{y})$ and $\text{Normal}(0,s_{y^{\prime}})$ respectively. This is a common technique used to obtain weakly-informative priors for scale parameters (Bürkner,, 2017). The prior on the length scale $\rho\sim\text{InvGamma}(5,5)$ is roughly based on the mean Euclidean distance between the $\tilde{x}_{i}$ (as in Section 3.3). As prior for the between-output correlation matrix $C$, we apply an unimodal LKJ(3) distribution (Lewandowski et al.,, 2009) defined over the positive definite symmetric matrices with unit diagonals. This distribution is a common prior choice for correlation matrices.
All models were specified in Stan (Stan Development Team,, 2024) and fitted with a single MCMC chain of 3000 iterations in total with 1000 warm-ups. We decided to run only a single chain per model to reduce overall computation times. All models were fitted on all generated datasets. The full study design is depicted in Figure 1.
@@ 4.3 Summary methods
where the expectation is taken over the posterior (approximated via samples). $\mathbb{E}(x_{post}-x_{true})^{2}$ can be decomposed into $\text{Var}(x_{post})$ and $\text{Bias}(x_{post},x_{true})^{2}$, thus measuring Bias-variance trade-off. We compute RMSE from all fitted models shown in Table 1 for each set of output dimensions (5, 10 and 20). We prefer models that provide both low bias indicating posterior mean estimates close to the ground truth as well as lower posterior variance indicating high precision, together resulting in an overall low RMSE.
To analyse RMSE values, we use a multilevel analysis of variance model (ANOVA) fitted with brms (Bürkner,, 2017), which disentangles the contributions of each model component. Using a multilevel model is important to account for the dependency between results of all models fitted on the same dataset. We model fixed main effects of scaled derivatives, varying hyperparameters, correlated outputs and number of output dimensions. For this purpose, we consider scaled derivatives as a factor variable with three levels corresponding to models that (a) do not include derivative information, (b) models that include derivative information with scaling and (c) models that include derivative information without scaling. Varying hyperparameters are represented by a binary factor variable that denotes varying vs. constant hyperparameters across output dimensions. Similarly, correlated outputs represented by a binary factor variable that indicates if the multiple outputs are assumed to be correlated or not. We additionally model fixed interaction effects between (a) scaled derivatives and varying hyperparameters and (b) scaled derivatives and correlated outputs. Since our simulation study is performed over 5, 10, and 20 output dimensions, we include dimension as factor variable with three levels and allowed it to interact with all previously mentioned (fixed) main and interaction effects. We account for the dependency structure in the RMSE values, induced by fitting multiple models to the same simulated dataset, by a random intercept over datasets as well as corresponding random slopes of the scaled derivatives, varying hyperparameters, and correlated outputs factors. Further, we account for the dependency in the RMSE values for the 20 latent inputs estimated from a single model through a random intercept per fitted model.
@@ 4.4 Model convergence
Figure 2: Validation measures for (a) latent inputs and (b) GP hyperparameters for the simulated GP scenario. The individual points correspond to each fitted models.
We investigate the convergence of our fitted GP models using standard MCMC sampling diagnostics including state-of-the art versions of the scale-reduction factor $\widehat{R}$, bulk effective sample size (Bulk-ESS) and tail effective sample size (Tail-ESS) (Vehtari et al.,, 2021). The combined check of these measures provides a comprehensive picture of individual parameters’ model convergence.
Figure 3: Validation measures for (a) latent inputs and (b) GP hyperparameters for the simulated periodic scenario. The individual points correspond to each fitted models.
In general, $\widehat{R}$ should be very close to 1 and should ideally not exceed 1.01 (Vehtari et al.,, 2021). In a simulation setup, we can evaluate the goodness of the posterior estimation also independently of convergence, as we have access to ground truth values. Hence, also in light of the relatively short MCMC chains, we decide to apply a more relaxed threshold of 1.1. Bulk-ESS indicates the reliability of measures of central tendency such as the posterior mean or median. Tail-ESS indicates the reliability of the 5% and 95% quantile estimates, which are commonly used to construct credible intervals. Both Bulk-ESS and Tail-ESS should have values higher than 100 times the number of MCMC chains. We computed all convergence measures with the posterior package (Bürkner et al.,, 2023).
For both hyperparameters and latent inputs obtained from the simulated GP scenario, we show $\widehat{R}$, Bulk-ESS, and Tail-ESS in Figure 2. The $\widehat{R}$ were all satisfactory with the exception of a few outlying models fitted on 20-dimensional outputs. Bulk-ESS and Tail-ESS were consistently higher than 100, thus satisfying the recommended criteria. Convergence results were similar for periodic scenario as shown in Figure 3.
@@ 4.5 Results
For both GP and periodic simulation scenarios, we evaluated the effects of including derivative information (Figure 4). In addition to the posterior RMSE estimates obtained from the multilevel ANOVA, we also show the prior RMSE that would be expected if we only used the prior measurement model $\tilde{x}_{i}\sim\mathcal{N}(x_{i},s^{2})$ to infer $x$. Accordingly, the prior RMSE acts as a benchmark to illustrate how much precision we gain through the GP modeling of output data. We see how the inclusion of both derivative information and scaling modifications simultaneously results in an overall substantial decrease in mean RMSE thus indicating a better recovery of the true latent values as compared to models without derivative information. Additionally, the mean RMSE further decreases as we increase the number of output dimensions.
Figure 4: Summarizing main effects of scaled derivatives for simulated (a) GP data and (b) periodic data.
Figure 5: Summarizing interaction effects of scaled derivatives with varying hyperparameters for simulated (a) GP data and (b) periodic data.
Figure 6: Summarizing interaction effects of scaled derivatives with correlated outputs for simulated (a) GP data and (b) periodic data.
@@
Overall, we see an RMSE reduction of more than 50% compared to the prior RMSE, and a reduction of about 30% compared to the models without derivative information, thus clearly outlining the benefits of using DGP-LVMs. Conversely, when models include derivative information without accounting for scale differences, RMSE values are a lot higher, suggesting that the model performs adversely while estimating latent inputs (Figure 11 in Appendix B). Curiously, the performance of such models is even worse than the models not including derivatives at all, sometimes even worse than when just using the prior measurement model alone. Presumably, this is because hyperparameter estimates are strongly biased if forced to be the same for both regular outputs and their derivatives; at least when the ground truth assumes hyperparameters to be different by a factor of 3 (which is not unrealistic). As an implication, we then also obtain strongly biased latent input estimates, resulting in large RMSEs. This clearly highlights the importance of our derivative covariance function modifications. Without it, using derivative information poses the risk of providing strongly misleading results.
Figure 7: Hyperparameter RMSEs with respect to their ground truth for the full DGP-LVM including all components. The different colour denotes if the hyperparameters correspond to the original or the derivative part of the model.
Figure 8: Hyperparameter RMSEs for models with scaled derivatives (a) without varying hyperparameters, (b) without correlated outputs and (c) without both varying hyperparameters and correlated outputs. The different colour denotes if the hyperparameters correspond to the original or the derivative part of the model.
@@
With respect to the other varied components, modeling varying hyperparameters results in a slight increase in RMSE, especially in higher output dimensions (Figure 5). We hypothesize that this is due to the significant increase in the number of estimated parameters, while the amount of data remain constant. Concretely, the number of parameters increase by the number of hyperparameters times the number of output dimensions, which is quite substantial especially for 20 output dimensions. It is likely that, in more data-rich scenarios, modeling hyperparameters as varying over dimensions is strongly beneficial. However, due to the increased computational costs of such a scenario, we do not investigate it in this paper. We see a similar increase in RMSE when modeling outputs as correlated, the reason presumably also being the substantial increase in estimated model parameters. For output dimension $D$, we estimate $D(D-1)/2$ number of parameters just for between-dimension correlations. Such a significant increase in parameters becomes visible in the results especially for $D=20$, as can be seen in Figure 6.

### v3

@@ 4.1 Simulated data
We consider five primary scenarios to generate simulated data. In our first scenario, we generate data from a multi-output GP with scaled derivative SE covariance function. In our second and third scenarios, we generate data from a multi-output GP but with scaled derivative Matern 3/2 and 5/2 covariance functions, respectively. All of the above scenarios constitute cases where the estimated DGP-LVMs align with the true underlying process. However, datasets generated this way can vary strongly in the amount of signal they contain, thus adding a lot of random variation in the simulation results. To account for this issue, in our fourth scenario, we generate data from a derivative periodic process with the true generating function
$f_{ij}$ $=\alpha_{j}\sin\left(\frac{x_{i}}{\rho_{j}}\right)$ (17)
$f^{\prime}_{ij}$ $=\frac{\alpha^{\prime}_{j}}{\rho_{j}}\cos\left(\frac{x_{i}}{\rho_{j}}\right)$
and corresponding data simulations as
@@
In all of these scenarios, the data is generated with varying hyperparameters and correlated outputs assumptions in play (see Section 3). The hyperparameters of the periodic data generating process fulfil a similar purpose to those of SE, Matern 3/2 and Matern 5/2 derivative GPs. Hence we choose to use the same hyperparameter names for simplicity. The scenario of periodic data (Eq.(17)) is important on two counts. First, it allows us to better control the amount of signal contained in each generated dataset. Second, it demonstrates that DGP-LVM can achieve good results even when the underlying generating process in not actually a GP.
Additionally, in the fifth scenario, we further increase the complexity by adding a quadratic and linear trend to the above periodic and corresponding derivative functions, respectively, as
$f_{ij}$ $=\alpha_{j}\sin\left(\frac{x_{i}}{\rho_{j}}\right)+bx_{i}^{2},$ (19)
$f^{\prime}_{ij}$ $=\frac{\alpha^{\prime}_{j}}{\rho_{j}}\cos\left(\frac{x_{i}}{\rho_{j}}\right)+2bx_{i},$
where $b$ is a scaler and is chosen in accordance with the acting periodicity parameter $\rho$. The data generating process then follows Eq. (18). This fifth scenario is designed to test the limitations of modelling non-stationary data with stationary GPs.
@@
To demonstrate the adversity of scale difference between $y$ and $y^{\prime}$, we induce a scaling factor of $\lambda=3$ that propagates through the GP marginal SD and error SD (see Section 3.1.1). The GP marginal SD for the output $y$ and the derivative $y^{\prime}$ are related through $\lambda$ such that $\alpha=\lambda\alpha^{\prime}$. Similarly for the error SD, $\sigma=\lambda\sigma^{\prime}$. Therefore, we only specify sampling distributions for $\alpha^{\prime}\sim\text{Normal}^{+}(3,0.25^{2})$ and $\sigma^{\prime}\sim\text{Normal}^{+}(1,0.25^{2})$. In reality, $\lambda>3$ or $\lambda<1/3$ may very well occur (see Section 5). In the simulations, we chose to avoid more extreme $\lambda$ values to prevent substantial convergence issues for models without scaling modifications. This allows us to showcase these models’ (reduced) performance without confounding this finding with convergence considerations. The ground truth GP length scale is sampled as $\rho\sim\text{Normal}^{+}(1,0.05^{2})$. The choice of our sampling distributions, especially for the length scale $\rho$ being an informed prior, enables us to explicitly select a range of values for our hyperparameters for which the simulated GP data would contain sufficient amount of signal. Further, we introduce an uniform between-dimension correlation of $0.5$ for all the simulated data scenario as to represent moderate interactions between outputs. Combined with the sampling of true hyperparameters for each output dimension, our simulated data mimics the real-world data scenario where such assumptions are prevalent.
@@
Lastly, we generate ground truth for $x$ as a sequence of values between $\{0.5,...,9.5\}$ with a total output sample size of $N=20$ for $y$ and $y^{\prime}$ each and we choose a value for prior measurement SD of the noisy $\tilde{x}$ as $s=0.3$ (see Section 3.3). This resembles the realistic scenario where observed times are already a relatively good measure of latent pseudotime considering the overall input scale. Both simulation studies are generated as multi-output data with three sets of output dimensions namely $D=2,5\text{ }\text{and}\text{ }10$. To keep model estimation times within manageable bounds, each simulated dataset contains only 20 $y$ and correspondingly 20 $y^{\prime}$ sample points. We perform 50 trials for each simulation scenario, that is, generate 50 datasets for the three GP data, the periodic and the periodic with trend scenario, respectively.
@@ 4.2 Model setup
The models within our DGP-LVM framework can vary specifically in four components, namely the inclusion of (1) derivative information, (2) scaled derivatives, (3) varying hyperparameters, as well as (4) correlations across output dimensions. In order to study the individual importance of the four components, we systematically enable/disable each of them and investigate the resulting models’ performances. The underlying data generating process contains all of the above components, so any model that only has a subset of components will be misspecified at least to some degree. In our simulations, component combinations were fully crossed where sensible (see Table 1 for an overview). We fit 12 GP models for each selected number of output dimensions $D$ resulting in a total of 36 models fitted per generated dataset for a specific simulation scenario. We only exclude specific, non-sensible combinations. For example, it does not make sense to ask if a GP model, which does not include derivative information, accounts for the scaling of the derivatives. We use a constant mean function (similar to an intercept in regression models) in all our models for both $f$ and $f^{\prime}$(wherever applicable) throughout the simulation study. Such specifications help with overall location shifts in the data.
@@
Prior specifications for all the model hyperparameters involved were aligned with the data generating conditions to a reasonable extent to better showcase the model contributions. We specify separate priors for the marginal SDs and error SDs corresponding to $f$, $f^{\prime}$ and $y$, $y^{\prime}$ respectively, to account for the scale differences between the original and derivative part of the data. We specify the priors for GP marginal SDs $\alpha\sim\text{Normal}^{+}(9,0.75^{2})$ and $\alpha^{\prime}\sim\text{Normal}^{+}(3,0.25^{2})$. In case of error SDs we specify $\sigma\sim\text{Normal}^{+}(3,0.75^{2})$ and $\sigma^{\prime}\sim\text{Normal}^{+}(1,0.25^{2})$. We shifted our priors for the original part of the GP $f$ and data $y$ in accordance with our choice of scale difference $\lambda$ in the simulation scenarios. We use an informative prior on the length scale $\rho\sim\text{Normal}^{+}(1,0.05^{2})$ is roughly based on the mean Euclidean distance between the $\tilde{x}_{i}$ (as in Section 3.3) as well as the data generating specifications showed in Section 4.1. As prior for the between-output correlation matrix $C$, we apply an unimodal LKJ(1) distribution (Lewandowski et al.,, 2009) defined over the positive definite symmetric matrices with unit diagonals. This distribution is a common prior choice for correlation matrices. For our constant mean function (intercepts), we used a Normal distribution with data-specific Mean and SD (for both $y$ and $y^{\prime}$ correspondingly) as prior.
All models were specified in Stan (Stan Development Team,, 2024) and fitted with a single MCMC chain of 3000 iterations in total with 1000 warm-ups. We decided to run only a single chain per model to reduce overall computation times. However, we show that using multiple chains yield similar results down the line (see Appendix: Figures 32 and 33.). All models were fitted on all generated datasets. The full study design is depicted in Figure 1.
@@
The simulation studies were conducted using 50 vCPUs (Intel(R) Xeon(R) Gold 6230R CPU @ 2.10 GHz) with 720 GB memory allowances. The maximum runtime (in hours) for DGP-LVM with $D=10$ (for a simulated dataset) were approximately 4.6 for the SE model, 5.9 for the Matern 3/2 model and 8.9 for the Matern 5/2 model.
@@ 4.3 Summary methods
and the mean absolute error:
$\text{MAE}(x_{post})=\mathbb{E}(\mid x_{post}-x_{true}\mid)$ (21)
where the expectations are taken over the posterior (approximated via samples). In case of RMSE, $\mathbb{E}(x_{post}-x_{true})^{2}$ can be decomposed into $\text{Var}(x_{post})$ and $\text{Bias}(x_{post},x_{true})^{2}$, thus measuring Bias-variance trade-off. We compute RMSE and MAE from all fitted models shown in Table 1 for each set of output dimensions (2,5 and 10). We prefer models that provide both low bias indicating posterior mean estimates close to the ground truth as well as lower posterior variance indicating high precision, together resulting in an overall low RMSE. Similarly, we prefer low MAE since it shows the amount of absolute bias present while estimating the latent variable. Overall, RMSE penalizes the models for estimating outlying posteriors while MAE is more lenient in that sense.
To analyse RMSE and MAE values, we use a multilevel analysis of variance model (ANOVA) fitted with brms (Bürkner,, 2017), which disentangles the contributions of each model component. Using a multilevel model is important to account for the dependency between results of all models fitted on the same dataset. We model fixed main effects of scaled derivatives, varying hyperparameters, correlated outputs and number of output dimensions. For this purpose, we consider scaled derivatives as a factor variable with three levels corresponding to models that (a) do not include derivative information, (b) models that include derivative information with scaling and (c) models that include derivative information without scaling. Varying hyperparameters are represented by a binary factor variable that denotes varying vs. constant hyperparameters across output dimensions. Similarly, correlated outputs represented by a binary factor variable that indicates if the multiple outputs are assumed to be correlated or not. We additionally model fixed interaction effects between (a) scaled derivatives and varying hyperparameters and (b) scaled derivatives and correlated outputs. Since our simulation study is performed over 2, 5, and 10 output dimensions, we include dimension as factor variable with three levels and allowed it to interact with all previously mentioned (fixed) main and interaction effects. We account for the dependency structure in the RMSE and MAE values, induced by fitting multiple models to the same simulated dataset, by a random intercept over datasets as well as corresponding random slopes of the scaled derivatives, varying hyperparameters, and correlated outputs factors. Further, we account for the dependency in the evaluation metric values for the 20 latent inputs estimated from a single model through a random intercept per fitted model. The results based on RMSE are presented in Section 4.5 while their corresponding MAE results are shown in Appendix B.
@@ 4.4 Model convergence
We investigate the convergence of our fitted GP models for all the five simulation scenarios mentioned before. To that end, we use standard MCMC sampling diagnostics including state-of-the-art versions of the scale reduction factor $\widehat{R}$, the bulk effective sample size (Bulk-ESS) and the tail effective sample size (Tail-ESS) (Vehtari et al.,, 2021). The combined check of these measures provide a comprehensive picture of individual parameter model convergence.
In general, $\widehat{R}$ should be very close to 1 and should ideally not exceed 1.01 (Vehtari et al.,, 2021). In a simulation setup, we can evaluate the goodness of the posterior estimation also independently of convergence, as we have access to ground truth values. Hence, also in light of the relatively short MCMC chains, we decide to apply a more relaxed threshold of 1.1. Bulk-ESS indicates the reliability of measures of central tendency such as the posterior mean or median. Tail-ESS indicates the reliability of the 5% and 95% quantile estimates, which are commonly used to construct credible intervals. Both Bulk-ESS and Tail-ESS should have values greater than 100 times the number of MCMC chains. We computed all the convergence measures with the posterior package (Bürkner et al.,, 2023).
Figure 2: Squared exponential scenario: Convergence measures for (a) latent inputs and (b) GP hyperparameters. The individual points correspond to each fitted models per simulated data. The y-axis for Bulk and Tail ESS plots are log10 transformed.
For latent inputs and hyperparameters obtained from the simulated SE data scenario, we show $\widehat{R}$, Bulk-ESS, and Tail-ESS in Figure 2. We present the MCMC diagnostics plots for the other simulated data scenarios (see Figures 14-17) in Appendix B owing to their similar nature. The $\widehat{R}$ were all satisfactory for majority of the simulation trials, with the exception of a few outlying models per trial. The GPs with derivative SE covariance function per simulation trial had better convergence as compared to the GPs with derivative Matern covariance functions. This was expected due to the increased model complexity of the Matern covariance functions. Moreover, the derivative Matern 3/2 (see Figure 14) being the boundary of existing derivative covariance functions among the Matern class makes it more complex for the sampler to perform as good as the Matern 5/2 (see Figure 15) and subsequently the much simpler SE covariance function. The Bulk-ESS and Tail-ESS were consistently higher than the suggested threshold for all the cases, thus satisfying the recommended criteria. The convergence results for the periodic and periodic with trend scenarios as shown in Figures 16 and 17 were similar to derivative SE data simulation scenario.
@@ 4.5 Results
For all of the simulation scenarios discussed in Section 4.1, we evaluated the effects of including derivative information, accounting for scale differences between $y$ and $y^{\prime}$, estimating varying hyperparameters across multiple outputs as well as correlated outputs. We summarize these aforementioned conditions as model assumptions and show their effects on the RMSE as our primary model evaluation measure of the posterior estimates of latent $x$ and covariance function hyperparameters with respect to their true simulated values. The corresponding MAE results were qualitatively highly similar and are thus only shown in Appendix B.
4.5.1 Model evaluation: Latent inputs
@@
In addition to the posterior model evaluation measure estimates obtained from multilevel ANOVA, we also show the prior RMSE that would be expected if we only used the prior measurement model $\tilde{x}_{i}\sim\mathcal{N}(x_{i},s^{2})$ to infer $x$. Consequently, the prior evaluation measure acts as a benchmark to illustrate how much precision we gain through the GP modelling of output data.
Figure 3: Squared exponential scenario: Main effects of including (a) scaled derivatives and interaction effects of assuming (b) varying hyperparameters and (c) correlated outputs on recovery of latent inputs
Our findings for the latent $x$ are presented for different simulation scenarios in Figures 35 for the simulated GP data scenarios and in Figure 6 for the periodic data scenario. We see how the inclusion of both derivative information and scaling modifications simultaneously results in an overall substantial decrease in mean RMSE in the simulated SE and periodic data scenarios (Figure 3(a) and 6(a)), thus indicating a better recovery of the true latent values as compared to models without derivative information. In case of the Matern 3/2 and 5/2 data scenarios, although not as substantial as the SE and periodic case, we see similar effects of adding scaled derivatives. This is due to the more challenging nature of the GPs with derivative Matern covariance functions as seen through their model convergence. Additionally, the evaluation measures for all the scenarios further decrease as we increase the number of output dimensions.
Figure 4: Matern 3/2 scenario: Main effects of including (a) scaled derivatives and interaction effects of assuming (b) varying hyperparameters and (c) correlated outputs on recovery of latent inputs
@@
Overall, we see a reduction in RMSE of more than 50% compared to the corresponding prior metrics, and a reduction of about 30% compared to the models without derivative information in the SE and periodic data cases, thus clearly outlining the benefits of using DGP-LVMs. Conversely, when models include derivative information without accounting for scale differences, the RMSEs are a lot higher, suggesting that the model performs adversely while estimating latent inputs (see Figures 1821 in Appendix B). Curiously, the performance of such models is even worse than the models not including derivatives at all, sometimes close to (or even worse than) when just using the prior measurement model alone. Presumably, this is because hyperparameter estimates are strongly biased if forced to be the same for both regular outputs and their derivatives; at least when the ground truth assumes hyperparameters to be different by a factor of 3 (which is not unrealistic). As an implication, we then also obtain strongly biased latent input estimates, resulting in large RMSEs. This clearly highlights the importance of our derivative covariance function modifications. Without these modifications, using derivative information poses the risk of providing strongly misleading results.
Figure 5: Matern 5/2 scenario: Main effects of including (a) scaled derivatives and interaction effects of assuming (b) varying hyperparameters and (c) correlated outputs on recovery of latent inputs
Figure 6: Periodic scenario: Main effects of including (a) scaled derivatives and interaction effects of assuming (b) varying hyperparameters and (c) correlated outputs on recovery of latent inputs
@@
With respect to the other varied components, modelling varying hyperparameters and correlated outputs may result in a slight increase in the RMSEs (see (b) and (c) of Figures 46), especially in higher output dimensions. We hypothesize that this is due to the significant increase in the number of estimated parameters, while the amount of data points remained constant in our simulations. Concretely, the number of parameters increase by the number of hyperparameters per dimension (i.e., $5$ in our case) times the number of output dimensions $D$, which is quite substantial already for $D=10$ output dimensions.
We encounter a similar issue when modelling outputs as correlated since the increase in estimated model parameters are even more substantial. For output dimension $D$, we estimate $D(D-1)/2$ number of parameters just for between-dimension correlations. Such a significant increase in parameters becomes visible in the results especially for $D=10$ for most of the simulated scenarios.
4.5.2 Model evaluation: Hyperparameters
Figure 7: Squared exponential scenario: Hyperparameter RMSEs for (a) full DGP-LVM model and (b) models without scale assumption. The different colour denotes if the hyperparameters correspond to the original or the derivative part of the model.
In Figures 710, we show hyperparameter recovery for the full DGP-LVM in the GP simulation scenarios and the periodic simulation scenario and compare the effects of accounting for scaling (each figure shows (a) full model with scaling and (b) without scaling).